write the equation of a line through (2,5) and (8,1) with it in slope intercept.
Found 2 solutions by bosun1, Jamaal:
Answer by bosun1(3)

write the equation of a line through (2,5) and (8,1) with it in slope intercept.
M= Y2-Y1/X2-X1 [FORMULAR]

NAME THE NUMBER'S FOR CONVINIENCY.
(2,5) and (8,1)
X1 Y1 X2 Y2


M= Y2-Y1/X2-X1 OR M=Y1-Y/X1-X
1-5 / 8-2 = -4/6
M=-4/6 .....-2/3 FINAL
FIND THE EQUATION OF THE LINE
FORMULA M=Y1-Y/X1-X

(2,5)
X1 Y1
-2/3=5-Y/2-X
CROSS MULTIPLY
GIVES YOU 2[5-Y]=-3[2-X]
10-2Y=-6+3X
-2Y=-6-10+3X
-2Y=-16+3X
DIVIDE BOTH SIDE BY -2
-2Y/-2=-16+3X/-2
Y=-16+3X/-2 FINAL ANSWER.

Answer by Jamaal(1)